# Is Ego Status All You Need for Open-Loop End-to-End Autonomous Driving?

### v1

@@ Appendix A Implementation Details
For the common practice, the future GT planning trajectory is generated from the ego locations of the samples in the subsequent 3 seconds. However, since one nuScenes clip is usually a 20s video, which means that the samples at the tail of the video (within 17s-20s) cannot produce a complete future trajectory, normal methods [16, 13] will perform special processing on these special samples by using masks, but ST-P3 [12] did not do this. ST-P3 mistakenly used samples from other scenes while generating GT of these tail samples, so errors occurred during training and testing.
Metrics Details.
@@
Ego Status Usage Details
Appendix B Post Optimization of UniAD.
@@
Appendix C Dropping Cameras
@@
Method Img Ego Status Intersetion $\downarrow$ Intersetion-ST (%) $\downarrow$ Intersetion-LR (%) $\downarrow$ Det. Map
@@
Table 2: Intersection-ST is the intersection rate with going straight driving commands. Intersection-LR is the intersection rate with turning left/right commands.

### v2

@@ Appendix A Implementation Details
For the common practice, the future GT planning trajectory is generated from the ego locations of the samples in the subsequent 3 seconds. However, since one nuScenes clip is usually a 20s video, which means that the samples at the tail of the video (within 17s-20s) cannot produce a complete future trajectory, normal methods [16, 13] will perform special processing on these special samples by using masks, but ST-P3 [12] did not do this. ST-P3 mistakenly used samples from other scenes while generating GT of these tail samples, so errors occurred during training and testing. Related issue: https://github.com/OpenDriveLab/ST-P3/issues/24.
Ego Status Usage Details
Appendix B Metrics Details.
Collision Rate.
$CR(t)=\frac{\sum_{i=0}^{N}\mathbb{I}_{i}}{N},N=t/0.5,$ (2)
$CR(t)={(\sum_{i=0}^{N}\mathbb{I}_{i}})>0,N=t/0.5.$ (3)
Method L2 (m) $\downarrow$ ($\sigma_{wd}$) $\downarrow$
1s 2s 3s Avg. 1s 2s 3s Avg.
Baseline 0.30 0.52 0.85 0.56 0.03 0.19 0.70 0.31
Table 1: The smoothness $\sigma_{wd}$ of predicted trajectories.
Trajectory Smoothness
We also assessed the stability of the model’s predicted trajectories. Given that the model predicts the trajectory for the next three seconds at each moment, it means that for every absolute moment in time $t$, the model predicted multiple waypoints at time $t$ from various preceding times. We see these different waypoints as a distribution. In non-extreme conditions, this distribution should be as concentrated as possible to ensure smoothness in the driving process, as shown in Fig. 2. To quantitatively analyze this distribution, we calculated the squared deviation distance of these distribution points, as shown in  Tab. 1. We found that this smoothness metric does not convey more information than the L2 metric, and we believe that this requires more exploration to verify the rationality of an metric.
Valid Samples
@@
Appendix C Neglected Ego Status in Perception Stage.
[…]
BEVFormer 41.6 51.7
BEVFormer 41.3 51.5
Appendix D Post Optimization of UniAD.
@@
Appendix E Dropping Cameras
@@
Method Img Ego Status CCR $\downarrow$ CCR-ST (%) $\downarrow$ CCR-LR (%) $\downarrow$ Det. Map
@@
Table 4: CCR-ST is the CCR rate with going straight driving commands. CCR-LR is the CCR rate with turning left/right commands.
@@
Failure Cases
Although the majority of scenarios in the nuScenes dataset are relatively straightforward, it does include certain challenging scenes, notably those involving continuous cornering. As shown in Fig. 6, we can observe that methods with various settings all yielded suboptimal predicted trajectories when navigating high-curvature bends. For challenging scenarios like cornering, where the system must continuously make evolving decisions, evaluating open-loop autonomous driving systems poses a significant challenge. One limitation of open-loop methods is that they do not suffer from cumulative errors. In detail, in the case of an extremely erroneous trajectory predicted at a given timestep, the trajectory starting point for the next timestep is still based on the GT trajectory. The metric we utilize, CCR, is adept at identifying low-quality trajectories. However, an appropriate metric that can effectively highlight high-quality trajectories remains an intriguing direction for further exploration.
t